Solve for x and y
24(y-1)+2(y-3)+y=4(x+2)+6

Answer by Math_prodigy(24)

24(y-1)+2(y-3)+y=4(x+2)+6
24y-24+2y-6+y=4x+8+6
27y-30=4x+14
solve for y:
27y=4x+44
y=(4x/27)+(44/27)
solve for x:
4x=27y-44
x=(27y/4)-11